Mary is a stock analyst. She has data for Stock ABC from July 1, 2008 to June 30, 2013 with price $32 and $37, respectively, and for Stock XYZ from October 1, 2012 to December 31, 2016 with price $126 and $222, respectively. Which stock has better historical performance?

a. stock ABC outperformed Stock XYZ

b. Stock ABC underperformed Stock XYZ

c. Stock ABC and Stock XYZ performed roughly the same

d. Lack of information to make judgement

If Mary decides to invest $10,000 on these two stocks with 1/3 on Stock ABC and the rest on Stock XYZ. How much would she expect to get if she holds her portfolio for two years? (round up)

a. 11032

b. 12740

c. 13792

d. 14580
Answer #1

Answer(1):

Firstly we need to calculate the returns on both the stocks:

Return on stock = (Closing value - Beginning value) / Beginning value

Return from stock ABC: (37-35) / 35 = 5.71%

Return from stock XYZ: (222-126) / 126 = 76.19%

So option "B" is correct that says, "Stock ABC under performed Stock XYZ"

Answer(2): She is expected to get-

Stock ABC: 10000*1/3 = 3000

3000*5.71% = $171.30

Stock XYZ: 10000 * 1/7 = 7000

7000 * 76.19% = $5333.30

Total dollar return = $5504.6

She expect to get if she holds her portfolio for two years: 5504.6 * 2 = $11009